A body is thrown horizontally and falls through a height of 80m while moving horizontally through a distance of 40m. With what initial velocity was the body thrown
1
Expert's answer
2018-10-17T15:42:09-0400

When the body is thrown horizontally

L=v_0 t

H=(gt^2)/2

So, initial velocity

v_0=L/t=L/√(2H/g)

=40/√((2×80 )/9.8)=9.9 m/s